OpenStudy (anonymous):

find the excluded values of the rational expression 3 _________________ n2 - 5n + 4

OpenStudy (anonymous):

excluded values of the rational expression is where n2 - 5n + 4=0 so find this 2 values of n by solving this equation

OpenStudy (anonymous):

\[5n^{3} + 4n^{2} -3 = 0\] \[n(5n^{2} + 4n) = 3\] \[n = 3 or 5n^{2} + 4n -3 = 0\]

OpenStudy (anonymous):

there is a 3 over the rational expression the answer can not be 0 as that is not one of the answers listed in the multiple choice

OpenStudy (anonymous):

Can you say what is the multiple choice?

OpenStudy (anonymous):

a). 4 and 1 b). -5 and 4 c). -4 and -1 d). 3 and -5

OpenStudy (anonymous):

I think it's 3 and -5 is right

OpenStudy (anonymous):

because I said n = 3 OR 5n2+4n−3=0

OpenStudy (anonymous):

oh I think you changed the problem

OpenStudy (anonymous):

It was n't like that

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Did you change the problem?

OpenStudy (anonymous):

No, its the same

OpenStudy (anonymous):

ok wait a second I will solve it

OpenStudy (anonymous):

No there is something wrong with numbers Can you make sure that you have the same problem?

OpenStudy (anonymous):

Ok I didn't put this right, n2 should be n^2

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

I know I mean the numbers? Is it right?

OpenStudy (anonymous):

Im not sure what your asking me

OpenStudy (anonymous):

Can you message me?

OpenStudy (anonymous):

What do you want me to message you?

OpenStudy (anonymous):

Can you write the problem again please?

OpenStudy (anonymous):

3 ______________ n^2 - 5n + 4

OpenStudy (anonymous):

it doesn't have any solution from the multiple choice you have

OpenStudy (anonymous):

by ^ i mean to the power of

OpenStudy (anonymous):

I know that

OpenStudy (anonymous):

and the full sentence is find the excluded values of the rational expression (then what i wrote)

OpenStudy (anonymous):

Your sure theres no answer?

OpenStudy (anonymous):

No wait, Now I know what you want to solve.

OpenStudy (anonymous):

to find the excluded values, you have to make n2-5n+4=0 and then solve this equation (n+1)(n+4) n= -1 OR n= -4 Do you have these answers in the multiple choice?

OpenStudy (anonymous):

yes i do, but can you show me how you got that answer?

OpenStudy (anonymous):

I told you To find the excluded values you have to have something like 1/0 or anything undefind So to make your equation undefined, you have to set n2-5n+4=0 to make it 3/0 right? then solve the equation to find values for n (n+4)(n+1) then, n= -4 Or n= -1

OpenStudy (anonymous):

Oh, ok. Thanks for your help!

OpenStudy (anonymous):

No problem ur welcome Do you have any other questions?

OpenStudy (anonymous):

No, thanks though.

OpenStudy (anonymous):

message me if you want me solve any problem OK?

OpenStudy (anonymous):

ok